OpenStudy (anonymous):

Simplify if possible: 3x + 7x - 10

OpenStudy (saifoo.khan):

x = 1

jhonyy9 (jhonyy9):

3x + 7x -10 = 10x - 10 = 10 (x-1)

OpenStudy (anonymous):

Thanks ! :)

jhonyy9 (jhonyy9):

please